OpenStudy (anonymous):

Find the general form of the equation of the line that passes through the points (2,-2) and (5,2)

OpenStudy (anonymous):

u know the standard form?

OpenStudy (anonymous):

@ArifL is it Ax+By=C

OpenStudy (anonymous):

there are another

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

m=slope b= y intercept

OpenStudy (anonymous):

u know how to find the slope?

OpenStudy (anonymous):

@ArifL oh yeah i know that way but i need it in the other form. I got 4/3 for the slope, but I'm not sure what to do after that.

OpenStudy (anonymous):

y-y1=m(x-x1)